You have been hired as a market analyst and your first task is to estimate the own-price elasticity for one of the products that your employer sells in a competitive market. In reviewing the historical data, you note that when the market price was $72.15 a unit, the firm sold 16,412 million units a month. When the firm decreased price to $37.85 a unit, the firm sold 26,980 units. Estimate the own-price elasticity of demand to the hundredths place (two decimal places).
Answer #1

% change in market price = ($37.85 - $72.15)/$72.15 * 100 = -34.3/72.15 * 100 = -0.47 * 100 = -47%

% change in quantity demanded = (26,980 - 16,412)/16,412 * 100 = 10,568/16,412 * 100 = 0.64 * 100 = 64%

Own price elasticity of demand = % change in quantity demanded/% change in price = 64%/-47% = -1.36

Answer = -1.36